# The theory of implicit operations

### v1

@@ 1. Formulas, compactness, and preservation theorems
The following may be regarded as a converse to 1.1(ii) under some additional assumptions.
@@
As $\varphi$ is preserved by homomorphisms in $\mathsf{K}$, 1.1(i) yields an existential positive formula $\psi$ that is equivalent in $\mathsf{K}$ to $\varphi$. Since $\psi$ is an existential positive formula, it is equivalent to a finite disjunction $\psi_{1}\sqcup\dots\sqcup\psi_{m}$ of pp formulas $\psi_{i}$. It is then sufficient to show that there exists $i\leqslant m$ such that $\mathsf{K}\vDash\psi\leftrightarrow\psi_{i}$. Since $\mathsf{K}\vDash\psi\rightarrow(\psi_{1}\sqcup\dots\sqcup\psi_{m})$, by 1.4 there exists $i\leqslant m$ such that $\mathsf{K}\vDash\psi\rightarrow\psi_{i}$. Because $\psi$ is equivalent to $\psi_{1}\sqcup\dots\sqcup\psi_{m}$, we have $\mathsf{K}\vDash\psi_{i}\rightarrow\psi$. Thus, the formula $\psi$, and hence also $\varphi$, is equivalent in $\mathsf{K}$ to $\psi_{i}$, which is a pp formula. ∎
@@ 2. Universal algebra
While every variety is a quasivariety and every quasivariety is a universal class, the converses are not true in general. We call proper the quasivarieties that are not varieties and the universal classes that are not quasivarieties. Examples of a proper quasivariety and a proper universal class are the classes of cancellative commutative monoids (8.8) and of fields (3.16), respectively.
@@
The following is a straightforward consequence of 2.1.
@@
The 2.9 readily implies that $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rsi}})$ for every quasivariety $\mathsf{K}$. It is well known that this result can be improved by restricting to the class $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$ of finitely generated members of $\mathsf{K}_{\textsc{rsi}}$. As we were unable to find a reference in the literature, we provide a proof.
@@
From [58, Prop. 2.1.18] it follows that $\mathsf{K}=\mathbb{Q}(\mathsf{K}^{\textup{fg}})$, where $\mathsf{K}^{\textup{fg}}$ is the class of the finitely generated members of $\mathsf{K}$. Then let $\bm{A}\in\mathsf{K}^{\textup{fg}}$. By the 2.9 there exists a subdirect embedding $f\colon\bm{A}\to\prod_{i\in I}\bm{A}_{i}$, where $\bm{A}_{i}\in\mathsf{K}_{\textsc{rsi}}$ for every $i\in I$. As $f$ is a subdirect embedding, each $\bm{A}_{i}$ is a homomorphic image of $\bm{A}$ and, therefore, finitely generated. So, $\bm{A}_{i}\in\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$ for every $i\in I$. Since $f$ is an embedding into a direct product of members of $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$, we obtain $\bm{A}\in\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Thus, $\mathsf{K}^{\textup{fg}}\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$, and hence $\mathbb{Q}(\mathsf{K}^{\textup{fg}})\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Together with $\mathsf{K}=\mathbb{Q}(\mathsf{K}^{\textup{fg}})$, this yields $\mathsf{K}\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Since $\mathsf{K}$ is a quasivariety containing $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$, we conclude that $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. ∎
@@ 3. Implicit operations
Then we proceed to prove the implication from right to left. Suppose that $f$ is defined by an existential positive formula $\varphi(x_{1},\dots,x_{n},y)$. To conclude that $f$ is an implicit operation of $\mathsf{K}$, it suffices to show that it is an operation of $\mathsf{K}$. Consider a homomorphism $h\colon\bm{A}\to\bm{B}$ with $\bm{A},\bm{B}\in\mathsf{K}$ and $a_{1},\dots,a_{n}\in A$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$. As $\varphi$ defines $f$, we have $\bm{A}\vDash\varphi(a_{1},\dots,a_{n},f^{\bm{A}}(a_{1},\dots,a_{n}))$. Since $\varphi$ is an existential positive formula, we can apply 1.1(i), obtaining $\bm{B}\vDash\varphi(h(a_{1}),\dots,h(a_{n}),h(f^{\bm{A}}(a_{1},\dots,a_{n})))$. Therefore, $\langle h(a_{1}),\dots,h(a_{n})\rangle\in\mathsf{dom}(f^{\bm{B}})$ and $h(f^{\bm{A}}(a_{1},\dots,a_{n}))=f^{\bm{B}}(h(a_{1}),\dots,h(a_{n}))$ because $\varphi$ defines $f$. Hence, we conclude that $f$ is an operation of $\mathsf{K}$. ∎
Let $f,f_{1},\dots,f_{n}$ be $m$-ary partial functions on a set $X$. We write $f=f_{1}\cup\dots\cup f_{n}$ to indicate that $f$ is the union of $f_{1},\dots,f_{n}$ when they are viewed as subsets of $X^{m}\times X$. This condition is equivalent to the requirements that $\mathsf{dom}(f)=\mathsf{dom}(f_{1})\cup\dots\cup\mathsf{dom}(f_{n})$ and that $f_{i}(x)=f(x)$ for all $i$ and $x\in\mathsf{dom}(f_{i})$. As a consequence of 3.9, we obtain the following.
@@
As $\mathsf{K}$ is an elementary class, we can apply 3.9, obtaining that $f$ is defined by an existential positive formula $\varphi$. We may assume that $\varphi=\varphi_{1}\sqcup\dots\sqcup\varphi_{n}$ for some pp formulas $\varphi_{1},\dots,\varphi_{n}$. Since $\varphi$ defines $f$, it is functional in $\mathsf{K}$. Together with $\varphi=\varphi_{1}\sqcup\dots\sqcup\varphi_{n}$, this implies that each $\varphi_{i}$ is functional in $\mathsf{K}$ and, therefore, defines a partial function $f_{i}$ on $\mathsf{K}$. As $\varphi_{i}$ is a pp formula, from Theorem 3.9 it follows that $f_{i}$ is an implicit operation of $\mathsf{K}$. Now, recall that $\varphi=\varphi_{1}\sqcup\dots\sqcup\varphi_{n}$ defines $f$ and $\varphi_{i}$ defines $f_{i}$ for each $i\leqslant n$. Thus, we conclude that $f^{\bm{A}}=f_{1}^{\bm{A}}\cup\dots\cup f_{n}^{\bm{A}}$ for each $\bm{A}\in\mathsf{K}$. ∎
In view of Corollary 3.10, implicit operations of elementary classes are obtained by gluing together implicit operations defined by pp formulas. This is the reason why the most fundamental implicit operations in mathematics are defined by pp formulas (as opposed to arbitrary existential positive formulas), as shown by the forthcoming examples. We denote by $\mathsf{imp}_{pp}(\mathsf{K})$ the set of implicit operations of a given class $\mathsf{K}$ that are defined by pp formulas and, when $\mathsf{K}=\{\bm{A}\}$, we often write $\mathsf{imp}_{pp}(\bm{A})$ instead of $\mathsf{imp}_{pp}(\mathsf{K})$.
@@
As the formula in the above display is a quasiequation for all $i,j\leqslant m$, 2.3(ii) implies that it is also valid in $\mathbb{Q}(\mathsf{K})$. Together with Claim 3.12, this guarantees that $\varphi$ is functional in $\mathbb{Q}(\mathsf{K})$. ∎
@@
Let $\mathsf{K}$ be a class of algebras. Then $\mathsf{imp}(\mathsf{K})$ and $\mathsf{imp}_{pp}(\mathsf{K})$ are closed under composition.
@@
Suppose now that in addition $g,f_{1},\dots,f_{n}\in\mathsf{imp}_{pp}(\mathsf{K})$. We can then assume that the formulas $\psi,\varphi_{1},\dots,\varphi_{n}$ that define $g,f_{1},\dots,f_{n}$ are pp formulas. Let $\chi$ be defined as in (8). Since $\psi$ and each $\varphi_{i}$ are pp formulas, they are of the form $\exists\vec{v}\psi^{\prime}$ and $\exists\vec{u}_{i}\varphi_{i}^{\prime}$, where $\vec{v}$ and $\vec{u}_{i}$ are finite sequences of variables and $\psi^{\prime}$ and $\varphi_{i}^{\prime}$ are finite conjunctions of equations. It is straightforward to verify that $\chi$ is equivalent to the formula obtained by pulling out the existential quantifiers $\exists\vec{v}$ and $\exists\vec{u}_{i}$ from the conjunction in (8), and hence $\chi$ is equivalent to a pp formula. Thus, $h$ is defined by a pp formula, and so $h\in\mathsf{imp}_{pp}(\mathsf{K})$. ∎
@@
We refer to $\varphi_{n}$ as to the $n$-th Isbell’s formula. Notice that Isbell’s formulas are pp formulas. It follows from [23, Lem. 4.4] that each Isbell’s formula is functional in the variety of monoids. Whence, from 3.11 we deduce the following.
@@
Isbell’s formulas and the implicit operations they define, which we term Isbell’s operations, will play a prominent role in the next sections (see 4.8, 4.9, and 14.4). ∎
@@
Recall from 3.17 that $\mathbb{Q}(\mathsf{Field})=\mathsf{RCRing}$. As $\varphi$ is a pp formula and is functional in $\mathsf{Field}$, we can apply 3.11, obtaining that $\varphi$ defines an implicit operation $f$ of $\mathsf{RCRing}$. Lastly, (10) ensures that $f^{\bm{A}}(a)=\mathsf{wi}(a)$ for every field $\bm{A}$ and $a\in A$. ∎
@@
in the language of rings. By Lagrange’s four squares theorem any nonnegative integer can be written as the sum of four integer squares (see, e.g., [3, Thm. 11-3]). Therefore, for all $a,b\in\mathbb{Z}$ we have that $\mathbb{Z}\vDash\varphi(a,b)$ if and only if $b\geqslant 0$ and $a^{2}=b^{2}$, which happens exactly when $b=\lvert a\rvert$. In particular, $\varphi$ is functional in $\mathbb{Z}$. Then 3.11 implies that $\varphi$ defines an implicit operation $f$ of the quasivariety $\mathsf{K}$ of rings generated by $\mathbb{Z}$ such that $f^{\mathbb{Z}}$ is the absolute value function.
@@ 4. Existential elimination
More precisely, recall from Corollary 3.10 that every implicit operation of $\mathsf{K}$ can be obtained by gluing together finitely many implicit operations defined by pp formulas. The main result of this section states that, if $\mathsf{K}$ has the amalgamation property, the study of its implicit operations can be further simplified by observing that each implicit operation defined by a pp formula is interpolated by one defined by a conjunction of equations (4.3). We term this phenomenon existential elimination because conjunctions of equations are obtained by removing existential quantifiers from pp formulas.
@@
A class of algebras $\mathsf{K}$ is said to have the amalgamation property when for every pair of embeddings $h_{1}\colon\bm{A}\to\bm{B}$ and $h_{2}\colon\bm{A}\to\bm{C}$ with $\bm{A},\bm{B},\bm{C}\in\mathsf{K}$ there exists a pair of embeddings $g_{1}\colon\bm{B}\to\bm{D}$ and $g_{2}\colon\bm{C}\to\bm{D}$ with $\bm{D}\in\mathsf{K}$ such that $g_{1}\circ h_{1}=g_{2}\circ h_{2}$.
@@
Given a class of algebras $\mathsf{K}$, we denote by $\mathsf{imp}_{eq}(\mathsf{K})$ the set of implicit operations of $\mathsf{K}$ defined by a conjunction of equations. When $\mathsf{K}=\{\bm{A}\}$, we often write $\mathsf{imp}_{eq}(\bm{A})$ instead of $\mathsf{imp}_{eq}(\mathsf{K})$. [^4] Although we will not rely on this fact, we will show in 4.16 that $\mathsf{imp}_{eq}(\mathsf{K})$ need not be closed under composition (cf. Proposition 3.13). The aim of this section is to establish the following interpolation result.
@@
    every member of $\mathsf{imp}_{pp}(\mathsf{K})$ can be interpolated by some member of $\mathsf{imp}_{eq}(\mathsf{K})$;
@@
    every member of $\mathsf{imp}(\mathsf{K})$ can be interpolated by a finite subset of $\mathsf{imp}_{eq}(\mathsf{K})$.
As the variety of monoids lacks the amalgamation property (see, e.g., [76, p. 100] [^5] As observed in [76, p. 108], the failure of the amalgamation property for the variety of monoids can be seen as a consequence of the corresponding result for the variety of semigroups proved in [75] (see also [37, Exa. 1]). ), it falls outside the scope of Theorem 4.3. This is reflected by the fact that this variety possesses implicit operations defined by pp formulas that cannot be interpolated by any implicit operation defined by a conjunction of equations, an example being every $n$-th Isbell’s operation for $n\geqslant 1$ (see 4.15). On the other hand, the variety of distributive lattices has the amalgamation property (see, e.g., [7, Thm. VII.8.4]) and, therefore, each of its implicit operations defined by pp formulas can be interpolated by one defined by a conjunction of equations.
@@
Both statements follow from 4.5: for (i) let $g\colon\bm{B}\to\bm{B}^{\prime}$ be the inclusion map, and for (ii) let $g\colon\bm{B}\to\bm{B}/\theta$ be the canonical surjection. ∎
@@
$\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=\{b\in B:$ $\text{ there exist }f\in\mathsf{imp}_{pp}(\mathsf{K})\text{ and }\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$
@@
As shown in the next result, the amalgamation property simplifies the task of describing dominions.
@@
Let $\mathsf{K}$ be a class of algebras closed under finite direct products with the amalgamation property and $\bm{A}\leqslant\bm{B}\leqslant\bm{C}$ with $\bm{B},\bm{C}\in\mathsf{K}$. Then
@@
We first prove the inclusion from left to right. As $\bm{B}\leqslant\bm{C}$, from 4.6(i) it follows that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{C})$. Since $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq B$ by definition, we obtain that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{C})\cap B$.
@@
Recall from the assumptions that $\bm{B}\leqslant\bm{C}$. Then let $i\colon\bm{B}\to\bm{C}$ be the inclusion map, which is always an embedding. As $i\colon\bm{B}\to\bm{C}$ and $f_{1}\colon\bm{B}\to\bm{D}$ are a pair of embeddings with $\bm{B},\bm{C},\bm{D}\in\mathsf{K}$, we can apply the assumption that $\mathsf{K}$ has the amalgamation property, obtaining a pair of embeddings $g_{1}\colon\bm{C}\to\bm{E}$ and $g_{2}\colon\bm{D}\to\bm{E}$ with $\bm{E}\in\mathsf{K}$ such that
@@
Since $i,g_{1},g_{2}$, and $f_{1}$ are embeddings, so are the compositions $g_{1}\circ i\colon\bm{B}\to\bm{E}$ and $g_{2}\circ f_{2}\colon\bm{B}\to\bm{E}$. Together with $\bm{E}\in\mathsf{K}$, this allows us to apply the amalgamation property of $\mathsf{K}$, obtaining a pair of embeddings $h_{1},h_{2}\colon\bm{E}\to\bm{F}$ with $\bm{F}\in\mathsf{K}$ such that

### v2

@@ 1. Formulas, compactness, and preservation theorems
The following may be regarded as a converse to Theorem 1.1(ii) under some additional assumptions.
@@
As $\varphi$ is preserved by homomorphisms in $\mathsf{K}$, Theorem 1.1(i) yields an existential positive formula $\psi$ that is equivalent in $\mathsf{K}$ to $\varphi$. Since $\psi$ is an existential positive formula, it is equivalent to a finite disjunction $\psi_{1}\sqcup\dots\sqcup\psi_{m}$ of pp formulas $\psi_{i}$. It is then sufficient to show that there exists $i\leqslant m$ such that $\mathsf{K}\vDash\psi\leftrightarrow\psi_{i}$. Since $\mathsf{K}\vDash\psi\rightarrow(\psi_{1}\sqcup\dots\sqcup\psi_{m})$, by Corollary 1.4 there exists $i\leqslant m$ such that $\mathsf{K}\vDash\psi\rightarrow\psi_{i}$. Because $\psi$ is equivalent to $\psi_{1}\sqcup\dots\sqcup\psi_{m}$, we have $\mathsf{K}\vDash\psi_{i}\rightarrow\psi$. Thus, the formula $\psi$, and hence also $\varphi$, is equivalent in $\mathsf{K}$ to $\psi_{i}$, which is a pp formula. ∎
@@ 2. Universal algebra
While every variety is a quasivariety and every quasivariety is a universal class, the converses are not true in general. We call proper the quasivarieties that are not varieties and the universal classes that are not quasivarieties. Examples of a proper quasivariety and a proper universal class are the classes of cancellative commutative monoids (Example 8.8) and of fields (Example 3.16), respectively.
@@
The following is a straightforward consequence of Theorem 2.1.
@@
The Subdirect Decomposition Theorem 2.9 readily implies that $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rsi}})$ for every quasivariety $\mathsf{K}$. It is well known that this result can be improved by restricting to the class $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$ of finitely generated members of $\mathsf{K}_{\textsc{rsi}}$. As we were unable to find a reference in the literature, we provide a proof.
@@
From [58, Prop. 2.1.18] it follows that $\mathsf{K}=\mathbb{Q}(\mathsf{K}^{\textup{fg}})$, where $\mathsf{K}^{\textup{fg}}$ is the class of the finitely generated members of $\mathsf{K}$. Then let $\bm{A}\in\mathsf{K}^{\textup{fg}}$. By the Subdirect Decomposition Theorem 2.9 there exists a subdirect embedding $f\colon\bm{A}\to\prod_{i\in I}\bm{A}_{i}$, where $\bm{A}_{i}\in\mathsf{K}_{\textsc{rsi}}$ for every $i\in I$. As $f$ is a subdirect embedding, each $\bm{A}_{i}$ is a homomorphic image of $\bm{A}$ and, therefore, finitely generated. So, $\bm{A}_{i}\in\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$ for every $i\in I$. Since $f$ is an embedding into a direct product of members of $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$, we obtain $\bm{A}\in\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Thus, $\mathsf{K}^{\textup{fg}}\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$, and hence $\mathbb{Q}(\mathsf{K}^{\textup{fg}})\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Together with $\mathsf{K}=\mathbb{Q}(\mathsf{K}^{\textup{fg}})$, this yields $\mathsf{K}\subseteq\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Since $\mathsf{K}$ is a quasivariety containing $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}$, we conclude that $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. ∎
@@ 3. Implicit operations
Then we proceed to prove the implication from right to left. Suppose that $f$ is defined by an existential positive formula $\varphi(x_{1},\dots,x_{n},y)$. To conclude that $f$ is an implicit operation of $\mathsf{K}$, it suffices to show that it is an operation of $\mathsf{K}$. Consider a homomorphism $h\colon\bm{A}\to\bm{B}$ with $\bm{A},\bm{B}\in\mathsf{K}$ and $a_{1},\dots,a_{n}\in A$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$. As $\varphi$ defines $f$, we have $\bm{A}\vDash\varphi(a_{1},\dots,a_{n},f^{\bm{A}}(a_{1},\dots,a_{n}))$. Since $\varphi$ is an existential positive formula, we can apply Theorem 1.1(i), obtaining $\bm{B}\vDash\varphi(h(a_{1}),\dots,h(a_{n}),h(f^{\bm{A}}(a_{1},\dots,a_{n})))$. Therefore, $\langle h(a_{1}),\dots,h(a_{n})\rangle\in\mathsf{dom}(f^{\bm{B}})$ and $h(f^{\bm{A}}(a_{1},\dots,a_{n}))=f^{\bm{B}}(h(a_{1}),\dots,h(a_{n}))$ because $\varphi$ defines $f$. Hence, we conclude that $f$ is an operation of $\mathsf{K}$. ∎
Let $f,f_{1},\dots,f_{n}$ be $m$-ary partial functions on a set $X$. We write $f=f_{1}\cup\dots\cup f_{n}$ to indicate that $f$ is the union of $f_{1},\dots,f_{n}$ when they are viewed as subsets of $X^{m}\times X$. This condition is equivalent to the requirements that $\mathsf{dom}(f)=\mathsf{dom}(f_{1})\cup\dots\cup\mathsf{dom}(f_{n})$ and that $f_{i}(x)=f(x)$ for all $i$ and $x\in\mathsf{dom}(f_{i})$. As a consequence of Theorem 3.9, we obtain the following.
@@
As $\mathsf{K}$ is an elementary class, we can apply Theorem 3.9, obtaining that $f$ is defined by an existential positive formula $\varphi$. We may assume that $\varphi=\varphi_{1}\sqcup\dots\sqcup\varphi_{n}$ for some pp formulas $\varphi_{1},\dots,\varphi_{n}$. Since $\varphi$ defines $f$, it is functional in $\mathsf{K}$. Together with $\varphi=\varphi_{1}\sqcup\dots\sqcup\varphi_{n}$, this implies that each $\varphi_{i}$ is functional in $\mathsf{K}$ and, therefore, defines a partial function $f_{i}$ on $\mathsf{K}$. As $\varphi_{i}$ is a pp formula, from Theorem 3.9 it follows that $f_{i}$ is an implicit operation of $\mathsf{K}$. Now, recall that $\varphi=\varphi_{1}\sqcup\dots\sqcup\varphi_{n}$ defines $f$ and $\varphi_{i}$ defines $f_{i}$ for each $i\leqslant n$. Thus, we conclude that $f^{\bm{A}}=f_{1}^{\bm{A}}\cup\dots\cup f_{n}^{\bm{A}}$ for each $\bm{A}\in\mathsf{K}$. ∎
In view of Corollary 3.10, implicit operations of elementary classes are obtained by gluing together implicit operations defined by pp formulas. This is the reason why the most fundamental implicit operations in mathematics are defined by pp formulas (as opposed to arbitrary existential positive formulas), as shown by the forthcoming examples. We denote by $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ the set of implicit operations of a given class $\mathsf{K}$ that are defined by pp formulas and, when $\mathsf{K}=\{\bm{A}\}$, we often write $\mathsf{imp}_{\textsc{pp}}(\bm{A})$ instead of $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$.
@@
As the formula in the above display is a quasiequation for all $i,j\leqslant m$, Corollary 2.3(ii) implies that it is also valid in $\mathbb{Q}(\mathsf{K})$. Together with Claim 3.12, this guarantees that $\varphi$ is functional in $\mathbb{Q}(\mathsf{K})$. ∎
@@
Let $\mathsf{K}$ be a class of algebras. Then $\mathsf{imp}(\mathsf{K})$ and $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ are closed under composition.
@@
Suppose now that in addition $g,f_{1},\dots,f_{n}\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. We can then assume that the formulas $\psi,\varphi_{1},\dots,\varphi_{n}$ that define $g,f_{1},\dots,f_{n}$ are pp formulas. Let $\chi$ be defined as in (8). Since $\psi$ and each $\varphi_{i}$ are pp formulas, they are of the form $\exists\vec{v}\psi^{\prime}$ and $\exists\vec{u}_{i}\varphi_{i}^{\prime}$, where $\vec{v}$ and $\vec{u}_{i}$ are finite sequences of variables and $\psi^{\prime}$ and $\varphi_{i}^{\prime}$ are finite conjunctions of equations. It is straightforward to verify that $\chi$ is equivalent to the formula obtained by pulling out the existential quantifiers $\exists\vec{v}$ and $\exists\vec{u}_{i}$ from the conjunction in (8), and hence $\chi$ is equivalent to a pp formula. Thus, $h$ is defined by a pp formula, and so $h\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. ∎
@@
We refer to $\varphi_{n}$ as to the $n$-th Isbell’s formula. Notice that Isbell’s formulas are pp formulas. It follows from [23, Lem. 4.4] that each Isbell’s formula is functional in the variety of monoids. Whence, from Corollary 3.11 we deduce the following.
@@
Isbell’s formulas and the implicit operations they define, which we term Isbell’s operations, will play a prominent role in the next sections (see Example 4.8, Isbell’s Zigzag Theorem 4.9, and Proposition 14.4). ∎
@@
Recall from Theorem 3.17 that $\mathbb{Q}(\mathsf{Field})=\mathsf{RCRing}$. As $\varphi$ is a pp formula and is functional in $\mathsf{Field}$, we can apply Corollary 3.11, obtaining that $\varphi$ defines an implicit operation $f$ of $\mathsf{RCRing}$. Lastly, (10) ensures that $f^{\bm{A}}(a)=\mathsf{wi}(a)$ for every field $\bm{A}$ and $a\in A$. ∎
@@
in the language of rings. By Lagrange’s four squares theorem any nonnegative integer can be written as the sum of four integer squares (see, e.g., [3, Thm. 11-3]). Therefore, for all $a,b\in\mathbb{Z}$ we have that $\mathbb{Z}\vDash\varphi(a,b)$ if and only if $b\geqslant 0$ and $a^{2}=b^{2}$, which happens exactly when $b=\lvert a\rvert$. In particular, $\varphi$ is functional in $\mathbb{Z}$. Then Corollary 3.11 implies that $\varphi$ defines an implicit operation $f$ of the quasivariety $\mathsf{K}$ of rings generated by $\mathbb{Z}$ such that $f^{\mathbb{Z}}$ is the absolute value function.
@@ 4. Existential elimination
More precisely, recall from Corollary 3.10 that every implicit operation of $\mathsf{K}$ can be obtained by gluing together finitely many implicit operations defined by pp formulas. The main result of this section states that, if $\mathsf{K}$ has the amalgamation property, the study of its implicit operations can be further simplified by observing that each implicit operation defined by a pp formula is interpolated by one defined by a conjunction of equations (Theorem 4.3). We term this phenomenon existential elimination because conjunctions of equations are obtained by removing existential quantifiers from pp formulas.
@@
Given a class $\mathsf{K}$ of similar algebras, we say that
  1. (i)
    a tuple $\langle\bm{A},\bm{B},\bm{C},h_{1},h_{2}\rangle$ is a span in $\mathsf{K}$ when $h_{1}\colon\bm{A}\to\bm{B}$ and $h_{2}\colon\bm{A}\to\bm{C}$ is a pair of embeddings with $\bm{A},\bm{B},\bm{C}\in\mathsf{K}$;
  2. (ii)
    a span $\langle\bm{A},\bm{B},\bm{C},h_{1},h_{2}\rangle$ in $\mathsf{K}$ has an amalgam in $\mathsf{K}$ when there exists a pair of embeddings $g_{1}\colon\bm{B}\to\bm{D}$ and $g_{2}\colon\bm{C}\to\bm{D}$ with $\bm{D}\in\mathsf{K}$ such that $g_{1}\circ h_{1}=g_{2}\circ h_{2}$;
@@
  3. (iii)
    a member $\bm{A}$ of $\mathsf{K}$ is an amalgamation base for $\mathsf{K}$ when every span in $\mathsf{K}$ of the form $\langle\bm{A},\bm{B},\bm{C},h_{1},h_{2}\rangle$ has an amalgam in $\mathsf{K}$;
  4. (iv)
    $\mathsf{K}$ has the amalgamation property when every span in $\mathsf{K}$ has an amalgam in $\mathsf{K}$.
@@
Given a class of algebras $\mathsf{K}$, we denote by $\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$ the set of implicit operations of $\mathsf{K}$ defined by a conjunction of equations. When $\mathsf{K}=\{\bm{A}\}$, we often write $\mathsf{imp}_{\textsc{eq}}(\bm{A})$ instead of $\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$. [^4] Although we will not rely on this fact, we will show in Example 4.16 that $\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$ need not be closed under composition (cf. Proposition 3.13). The aim of this section is to establish the following interpolation result.
@@
    every member of $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ can be interpolated by some member of $\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$;
@@
    every member of $\mathsf{imp}(\mathsf{K})$ can be interpolated by a finite subset of $\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$.
As the variety of monoids lacks the amalgamation property (see, e.g., [76, p. 100] [^5] As observed in [76, p. 108], the failure of the amalgamation property for the variety of monoids can be seen as a consequence of the corresponding result for the variety of semigroups proved in [75] (see also [37, Exa. 1]). ), it falls outside the scope of Theorem 4.3. This is reflected by the fact that this variety possesses implicit operations defined by pp formulas that cannot be interpolated by any implicit operation defined by a conjunction of equations, an example being every $n$-th Isbell’s operation for $n\geqslant 1$ (see Example 4.15). On the other hand, the variety of distributive lattices has the amalgamation property (see, e.g., [7, Thm. VII.8.4]) and, therefore, each of its implicit operations defined by pp formulas can be interpolated by one defined by a conjunction of equations.
@@
Both statements follow from Proposition 4.5: for (i) let $g\colon\bm{B}\to\bm{B}^{\prime}$ be the inclusion map, and for (ii) let $g\colon\bm{B}\to\bm{B}/\theta$ be the canonical surjection. ∎
@@
$\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=\{b\in B:$ $\text{ there exist }f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})\text{ and }\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$
@@
As shown in the next result, dominions in amalgamation bases are especially well behaved.
@@
Let $\mathsf{K}$ be a class of algebras closed under finite direct products and $\bm{A}\leqslant\bm{B}\leqslant\bm{C}$ with $\bm{B},\bm{C}\in\mathsf{K}$. If $\bm{B}$ is an amalgamation base for $\mathsf{K}$, then
@@
We first prove the inclusion from left to right. As $\bm{B}\leqslant\bm{C}$, from Corollary 4.6(i) it follows that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{C})$. Since $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq B$ by definition, we obtain that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{C})\cap B$.
@@
Recall from the assumptions that $\bm{B}\leqslant\bm{C}$. Then let $i\colon\bm{B}\to\bm{C}$ be the inclusion map, which is always an embedding. As $i\colon\bm{B}\to\bm{C}$ and $f_{1}\colon\bm{B}\to\bm{D}$ are a pair of embeddings with $\bm{C},\bm{D}\in\mathsf{K}$, we can apply the assumption that $\bm{B}$ is an amalgamation base for $\mathsf{K}$, obtaining a pair of embeddings $g_{1}\colon\bm{C}\to\bm{E}$ and $g_{2}\colon\bm{D}\to\bm{E}$ with $\bm{E}\in\mathsf{K}$ such that
@@
Since $i,g_{1},g_{2}$, and $f_{2}$ are embeddings, so are the compositions $g_{1}\circ i\colon\bm{B}\to\bm{E}$ and $g_{2}\circ f_{2}\colon\bm{B}\to\bm{E}$. Together with $\bm{E}\in\mathsf{K}$, we obtain another span $\langle\bm{B},\bm{E},\bm{E},g_{1}\circ i,g_{2}\circ f_{2}\rangle$ in $\mathsf{K}$. The assumption that $\bm{B}$ is an amalgamation base for $\mathsf{K}$ yields a pair of embeddings $h_{1},h_{2}\colon\bm{E}\to\bm{F}$ with $\bm{F}\in\mathsf{K}$ such that